OpenStudy (anonymous):

for WorldBoy: f(x)=-7/2x -3 g(x)=-2x+6/7

OpenStudy (anonymous):

ok, you understand the problem?

OpenStudy (anonymous):

WB, I'll assume the x is not in the denominator

OpenStudy (anonymous):

like amistre64 said, to do compositions, put f(x) or g(x) in for x

OpenStudy (anonymous):

no

OpenStudy (anonymous):

in this one, you cross multi

OpenStudy (anonymous):

f(x)=-7/2x -3 g(x)=-2x+6/7 f(g(x)) = -7/2g(x) -3 f(g(x)) = -7/2(-2x+6/7 ) -3 It's all about parentheses

OpenStudy (anonymous):

oh my, god this was the simples thing ever,

OpenStudy (anonymous):

thx, i won't need more help tonks.

OpenStudy (anonymous):

f(g(x)) = -7/2(-2x+6/7 ) -3 = 7x - 3 -3 when you distribute = 7x -6 when you combine like terms

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

bye